# Information geometry approach to quantum stochastic thermodynamics

### v1

@@ V.0.2 Quantum generalisation
where $j_{k}^{xy}(t)$ is the probability current between instantaneous eigenstates |$\ket{y(t)}$ and |$\ket{x(t)}$, and $f_{k}^{xy}(t)$, like in the classical case, can be interpreted as a generalised thermodynamic force. In addition, $\phi_{k}^{xy}(t)$ is the entropy flow associated with jump $\hat{L}_{k}$ at time $t$ resulting in a probability current $j^{xy}_{k}(t)$ from state |$\ket{y(t)}$ to |$\ket{x(t)}$.
The rate equation for $p_{x}(t)$ may then be rewritten in terms of the probability currents between instantaneous eigenstates
@@
The rate of the von Neumann is therefore given by
@@ VI Geometric bound on the von Neumann entropy rate
In this section, we will first demonstrate that the classical uncertainty relation between the geometric uncertainty in the path in state space and the time-integrated rate of information change, established in [35] for classical systems, remains valid for quantum dynamics.
We then make use of the decomposition of any QFI into a common metric-independent incoherent contribution and a metric-dependent coherent contribution, revised in section IV. The incoherent part is inherently linked to the system’s entropy dynamics as shown in section V.0.2. This connection enables us to extend a classical bound on the change in the entropy rate, as described in [35], to the quantum domain.
@@
$\delta=4\frac{\mathcal{J}-\mathcal{L}^{2}}{\tau^{2}}=\mathrm{E}[F_{Q}]-\mathrm{E}[\sqrt{F_{Q}}]^{2}\geq 0.$ (53)
The equation above captures the cumulative path deviation from the geodesic, which represents the trajectory of constant curvature and speed between the initial and final states. Since $\mathcal{J}-\mathcal{L}^{2}$ vanishes along the geodesic, where the QFI remains constant by definition, any non-zero difference indicates a deviation from this optimal path. Because of the positivity of $\mathcal{L}^{2}$ and $\delta$, one finds that
$\delta\leq\frac{8\mathcal{J}}{\tau^{2}}.$ (54)
Here it is helpful to rewrite the above inequality in terms of the time averaged Fisher information, which represents the time-averaged rate of information change, $\mathcal{I}=\mathrm{E}[F_{Q}]=\frac{1}{\tau}\int_{0}^{\tau}\mathrm{d}tF_{Q}$. The inequality then becomes
@@
Building on the above uncertainty relation, we decompose the QFI into its incoherent and coherent components, $F_{Q}=F_{Q}^{IC}+F_{Q}^{C}$ and $F_{Q}^{IC}=\mathcal{B}-\ddot{S}$. By performing the integral $\int_{0}^{\tau}F_{Q}\mathrm{d}t=\int_{0}^{\tau}\left(\mathcal{B}(t)+F_{Q}^{C}(t)-\ddot{S}(t)\right)\mathrm{d}t$, we find that the difference in the entropy rate between the final and initial states is bounded from above
$\begin{split}\Delta\dot{S}&\leq\mathcal{C}-\frac{\delta\tau}{2}+\int_{0}^{\tau}\mathrm{d}tF_{Q}^{C},\\ \end{split}$ (56)
where $\Delta\dot{S}=\dot{S}(\tau)-\dot{S}(0)$ and $\mathcal{C}=\int_{0}^{\tau}\mathcal{B}(t)\mathrm{d}t$, which is the integrated difference between the the average local information rate and the bulk information rate.
@@
Note that if one considers a driving protocol where the initial state is the fixed point of the initial instantaneous generator of the dynamics, such that $\dot{S}(0)=0$, the above serves as a bound on the instantaneous rate of entropy change $\dot{S}(\tau)$ at any later time $\tau$. We can further exploit the relationship between the incoherent contribution to the QFI and the entropic acceleration in Markovian open system dynamics obeying local detailed balance. In cases where the heat flow between the system and its environment vanishes, the resulting bound applies to the entropy production. Conversely, if the entropy production is zero, the bound constrains the entropic flow. It is also worth noting that our bound applies irrespective of the chosen QFI metric. However, it can be shown that $-\frac{\delta\tau}{2}+\int_{0}^{\tau}\mathrm{d}tF_{Q}^{C}$ is minimized for the SLD QFI, resulting in the tightest bound.
@@ VII Quantum Mpemba effect
We now apply the information-geometric framework developed in the previous sections to gain insight into the (thermo-)dynamics underlying a quantum Mpemba effect manifest during the thermal relaxation process of a qubit interacting with a thermal bosonic bath, as recently explored in [50]. Historically, the Mpemba effect [74, 75] describes the following phenomenon: consider two identical buckets of liquid. Both are at thermal equilibrium, but with respect to environments at different temperatures, $T_{h}$ and $T_{c}$, with $T_{h}>T_{c}$. Then, these buckets are placed in a third environment with a lower temperature $T_{h}>T_{c}>T^{*}$, and are allowed to thermalise. A Mpemba effect is said to be observed if there exists a time $t_{M}$ after which the temperature of the initially hotter liquid is lower than that of the initially colder liquid at all later times. The effect has stimulated a lot of theoretical and experimental interest [76, 77, 78, 79, 80, 81], however, a (universal) physical origin remains elusive and consensus on whether the effect exists in the first place has not yet been reached [82, 83].
More recently, Mpemba-like effects have been predicted and experimentally observed also in the quantum domain [84, 85, 86, 87, 88], primarily focused on symmetry-restoration dynamics in closed quantum systems.
@@
with equilibrium free energy $F_{\mathrm{eq}}=-\beta^{-1}\log Z_{\beta}$ and partition function $Z_{\beta}=\mathrm{Tr}\left[\exp(-\beta H)\right]$ and $\mathrm{D}(\hat{\sigma}\|\hat{\tau}_{\beta})$ is the quantum relative entropy between the relaxing state and the thermal fixed point. The authors in [50] defined a Mpemba effect to occur when a $\hat{\rho}_{0}^{\prime}$ with higher initial non-equilibrium free energy $F_{\mathrm{neq}}$ compared to the reference state $\hat{\rho}_{0}$, relaxes to thermal equilibrium faster. In particular, there exists a time $t_{M}$ after which $F_{\mathrm{neq}}(\hat{\rho}^{\prime}(t))>F_{\mathrm{neq}}(\hat{\rho}(t))$ for all $t>t_{M}$, indicating a crossing of non-equilibrium free energies along their respective relaxation paths.
@@
The qubit is initialised to a random mixed state, serving as the reference state $\hat{\rho}_{0}$. It is parametrised as $\hat{\rho}_{0}=\frac{1}{2}\left(1+\vec{r}\cdot\vec{\sigma}\right)$, where $\vec{\sigma}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})^{\mathrm{T}}$ and $\vec{r}=(r_{x},r_{y},r_{z})^{\mathrm{T}}$. We restrict to reference states with coherences in the initial energy eigenbasis so that at least $r_{x}$ or $r_{y}$ are non-zero. The unitarily rotated state, diagonal in the energy eigenbasis, is $\hat{\rho}^{\prime}_{0}=\mathrm{diag}(\lambda_{1},\lambda_{2})$, where $\lambda_{1}$ and $\lambda_{2}$ are the eigenvalues of $\hat{\rho}$, so that $\lambda_{1}\geq\lambda_{2}$. Importantly, the choice of $\hat{\rho}_{0}^{\prime}$ ensures that $F_{\mathrm{neq}}(\hat{\rho}_{0}^{\prime})\geq F_{\mathrm{neq}}(\hat{\rho}_{0})$.
The numerical results, which we will describe shortly, are obtained for the reference state and the rotated state, specified by the vectors $\vec{r}=(-0.41760,-0.60647,0.47879)$ and $\vec{r}^{\prime}=(0,0,0.87836)$, respectively. Additional parameters that determine the set-up are the temperature $T=10$, the Hamiltonian level splitting $\omega=5$ and the coupling strength to the bosonic reservoir $\gamma=1$.
Note that, in this context, the interaction with the environment cannot generate coherence in the energy eigenbasis. Thus, under dynamics induced by the GKSL master equation (Eq. 58), the rotated state will remain diagonal in the eigenbasis of the Hamiltonian at all later times. We now remember that a general time evolution can be understood as the combination of time-evolving eigenvalues, represented by $\chi(t)$ in a time evolving eigenbasis, expressed in $U_{t}$ (see section IV). But since the eigenbasis remains static, $U_{t}$ is the identity, and dynamics is captured fully by the time-evolving eigenvalues. Therefore, the QFI with respect to time reduces to its incoherent contribution - the classical Fisher information for the probability distribution constructed from the time-evolving eigenvalues. The reference state, however, initially has coherence in the energy eigenbasis by construction, and therefore the QFI has both an incoherent and a coherent contribution, as shown in Fig. 1b).
@@
We now shift focus to a geometric bound on the time-evolution of observables and their fluctuations. The bound we derived in section VI illustrates the possible benefit of choosing one QFI over another. Another advantage of the SLD is its direct link to the time-derivative of the density matrix. An interesting geometric consequence of this connection was established in [38]
@@ VIII Discussion
Finally, we show that the thermodynamic quantum Mpemba effect, recently reported in [50], can be understood through the analysis of geometric quantities, such as the ratio of completion and deviations from the geodesic path. Given the inherently general nature of the information-geometric framework, and its demonstrated applicability to quantum dynamics — while maintaining connections to thermodynamic quantities as shown in our work and others [38, 52] — we aim to use it in future research to uncover common underlying mechanisms driving various anomalous dynamics, such as Mpemba-like effects, in both closed and open quantum systems.

### v2

@@ V.0.2 Quantum generalisation
where $j_{k}^{xy}(t)$ is the probability current between instantaneous eigenstates |$\ket{y(t)}$ and |$\ket{x(t)}$, and $f_{k}^{xy}(t)$, like in the classical case, can be interpreted as a generalised thermodynamic force. In addition, $\phi_{k}^{xy}(t)$ is the entropy flow associated with the jump $\hat{L}_{k}$ at time $t$ resulting in a probability current $j^{xy}_{k}(t)$ from state |$\ket{y(t)}$ to |$\ket{x(t)}$. The rate equation for $p_{x}(t)$ may then be rewritten in terms of the probability currents between instantaneous eigenstates
@@
The rate of the von Neumann entropy is therefore given by
@@ VI Geometric bound on the von Neumann entropy rate
In this section, we will first demonstrate that the classical uncertainty relation between the geometric uncertainty in the path in state space and the time-averaged rate of information change, established in [35] for classical systems, remains valid for quantum dynamics. We then make use of the decomposition of any QFI into a common metric-independent incoherent contribution and a metric-dependent coherent contribution, revised in section IV. The incoherent part is inherently linked to the system’s entropy dynamics as shown in section V.0.2. This connection enables us to extend a classical bound on the change in the entropy rate, as described in [35], to the quantum domain.
@@
$\delta=4\frac{\mathcal{J}-\mathcal{L}^{2}}{T^{2}}=\mathrm{E}[F_{Q}]-\mathrm{E}[\sqrt{F_{Q}}]^{2}\geq 0,$ (53)
at any time $T$, so that $0\leq T\leq\tau$. The equation above captures the cumulative path deviation from the geodesic, which represents the trajectory of constant speed between the initial and final states. Since $\mathcal{J}-\mathcal{L}^{2}$ vanishes along the geodesic, where the QFI remains constant by definition, any non-zero difference indicates a deviation from this optimal path. Because of the positivity of $\mathcal{L}^{2}$ and $\delta$, one finds that
$\delta\leq\frac{8\mathcal{J}}{T^{2}}.$ (54)
Here it is helpful to rewrite the above inequality in terms of the time averaged Fisher information, which represents the time-averaged rate of information change, $\mathcal{I}=\mathrm{E}[F_{Q}]=\frac{1}{T}\int_{0}^{T}\mathrm{d}tF_{Q}$. The inequality then becomes
@@
Building on the above uncertainty relation, we decompose the QFI into its incoherent and coherent components, $F_{Q}=F_{Q}^{IC}+F_{Q}^{C}$ and $F_{Q}^{IC}=\mathcal{B}-\ddot{S}$. By performing the integral $\int_{0}^{T}F_{Q}\mathrm{d}t=\int_{0}^{T}\left(\mathcal{B}(t)+F_{Q}^{C}(t)-\ddot{S}(t)\right)\mathrm{d}t$, we find that the difference in the entropy rate between the final and initial states is bounded from above
$\begin{split}\Delta\dot{S}&\leq\mathcal{C}-\frac{\delta T}{2}+\int_{0}^{T}\mathrm{d}tF_{Q}^{C},\\ \end{split}$ (56)
where $\Delta\dot{S}=\dot{S}(T)-\dot{S}(0)$ and $\mathcal{C}=\int_{0}^{T}\mathcal{B}(t)\mathrm{d}t$, which is the integrated difference between the the average local information rate and the bulk information rate, for $0\leq T\leq\tau$.
@@
Note that if one considers a driving protocol where the initial state is the fixed point of the initial instantaneous generator of the dynamics, such that $\dot{S}(0)=0$, the above serves as a bound on the instantaneous rate of entropy change $\dot{S}(T)$ at any later time $T$, which indicates that the state is out of equilibrium. We can further exploit the relationship between the incoherent contribution to the QFI and the entropic acceleration in Markovian open system dynamics obeying local detailed balance. In cases where the heat flow between the system and its environment vanishes, the resulting bound applies to the entropy production. Conversely, if the entropy production is zero, the bound constrains the entropic flow. It is also worth noting that our bound applies irrespective of the chosen QFI metric. However, it can be shown that $-\frac{\delta T}{2}+\int_{0}^{T}\mathrm{d}tF_{Q}^{C}$ is minimized for the SLD QFI, resulting in the tightest bound.
@@ VII Quantum Mpemba effect
We now apply the information-geometric framework developed in the previous sections to gain insight into the (thermo-)dynamics underlying a quantum Mpemba effect manifest during the thermal relaxation process of a qubit interacting with a thermal bosonic bath, as recently explored in [51]. Historically, the Mpemba effect [75, 76] describes the following phenomenon: consider two identical buckets of liquid. Both are at thermal equilibrium, but with respect to environments at different temperatures, $T_{h}$ and $T_{c}$, with $T_{h}>T_{c}$. Then, these buckets are placed in a third environment with a lower temperature $T_{h}>T_{c}>T^{*}$, and are allowed to thermalise. A Mpemba effect is said to be observed if there exists a time $t_{M}$ after which the temperature of the initially hotter liquid is lower than that of the initially colder liquid at all later times. The effect has stimulated a lot of theoretical and experimental interest [77, 78, 79, 80, 81, 82], however, a (universal) physical origin remains elusive and consensus on whether the effect exists in the first place has not yet been reached [83, 84]. More recently, Mpemba-like effects have been predicted and experimentally observed also in the quantum domain [85, 86, 87, 88, 89], primarily focused on symmetry-restoration dynamics in closed quantum systems.
@@
with equilibrium free energy $F_{\mathrm{eq}}=-\beta^{-1}\log Z_{\beta}$, partition function $Z_{\beta}=\mathrm{Tr}\left[\exp(-\beta H)\right]$ and $\mathrm{D}(\hat{\sigma}\|\hat{\tau}_{\beta})$ is the quantum relative entropy between the relaxing state and the thermal fixed point. The authors in [51] defined a Mpemba effect to occur when a $\hat{\rho}_{0}^{\prime}$ with higher initial non-equilibrium free energy $F_{\mathrm{neq}}$ compared to the reference state $\hat{\rho}_{0}$, relaxes to thermal equilibrium faster. In particular, there exists a time $t_{M}$ after which $F_{\mathrm{neq}}(\hat{\rho}^{\prime}(t))>F_{\mathrm{neq}}(\hat{\rho}(t))$ for all $t>t_{M}$, indicating a crossing of non-equilibrium free energies along their respective relaxation paths.
@@
The qubit is initialised to an arbitrary mixed state, serving as the reference state $\hat{\rho}_{0}$. It is parameterised as $\hat{\rho}_{0}=\frac{1}{2}\left(1+\vec{r}\cdot\vec{\sigma}\right)$, where $\vec{\sigma}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})^{\mathrm{T}}$ and $\vec{r}=(r_{x},r_{y},r_{z})^{\mathrm{T}}$. We restrict to reference states with coherences in the initial energy eigenbasis so that at least $r_{x}$ or $r_{y}$ are non-zero. The unitarily rotated state, diagonal in the energy eigenbasis, is $\hat{\rho}^{\prime}_{0}=\mathrm{diag}(\lambda_{1},\lambda_{2})$, where $\lambda_{1}$ and $\lambda_{2}$ are the eigenvalues of $\hat{\rho}$, so that $\lambda_{1}\geq\lambda_{2}$. Importantly, the choice of $\hat{\rho}_{0}^{\prime}$ ensures that $F_{\mathrm{neq}}(\hat{\rho}_{0}^{\prime})\geq F_{\mathrm{neq}}(\hat{\rho}_{0})$.
The numerical results, which we will describe shortly, are obtained for the reference state and the rotated state, specified by the vectors $\vec{r}=(-0.41760,-0.60647,0.47879)$ and $\vec{r}^{\prime}=(0,0,0.87836)$, respectively. Additional parameters (with $\hbar=k_{\mathrm{B}}=1$) that determine the set-up are the coupling strength to the bosonic reservoir $\gamma=1$, the temperature $T=10$ and the Hamiltonian level splitting $\omega=5$.
Note that, in this context, the interaction with the environment cannot generate coherence in the energy eigenbasis. Thus, under dynamics induced by the GKSL master equation (Eq. 58), the rotated state will remain diagonal in the eigenbasis of the Hamiltonian at all later times. We now remember that a general time evolution can be understood as the combination of time-evolving eigenvalues, represented by $\chi(t)$ in a time evolving eigenbasis, expressed in $U_{t}$ (see section IV). But since the eigenbasis remains static, $U_{t}$ is the identity, and the dynamics are fully captured by the time-evolving eigenvalues. Therefore, the QFI with respect to time reduces to its incoherent contribution - the classical Fisher information for the probability distribution constructed from the time-evolving eigenvalues. The reference state, however, initially has coherence in the energy eigenbasis by construction, and therefore the QFI has both an incoherent and a coherent contribution, as shown in Fig. 1b).
@@
We now shift focus to a geometric bound on the time-evolution of observables and their fluctuations. The bound we derived in section VI illustrates the possible benefit of choosing one QFI over another. Another advantage of the SLD is its direct link to the time-derivative of the density matrix. An interesting geometric consequence of this connection for arbitrary observables $\hat{O}$ was first established in [38] (recently generalised also to other quantum Fisher metrics in [40]),
@@ VIII Conclusion
Finally, we show that the thermodynamic quantum Mpemba effect, recently reported in [51], can be understood through the analysis of geometric quantities, such as the ratio of completion and deviations from the geodesic path. Given the inherently general nature of the information-geometric framework, and its demonstrated applicability to quantum dynamics — while maintaining connections to thermodynamic quantities as shown in our work and others [38, 53, 40] — we aim to use it in future research to uncover common underlying mechanisms driving various anomalous dynamics, such as Mpemba-like effects, in both closed and open quantum systems.